# Prediction of linear fractional stable motions using codifference

### v1

@@ Appendix B Proof of Theorem 1
The autocodifference of the LFSM is provided in equation (4). But the knowledge of the symmetric codifference matrix of $(X_{t},...,X_{t+d-1})^{\prime}$ is equivalent to the knowledge of $\|X_{t+i}\|_{\alpha}$, for $10\leq i\leq d-$ 1, and $\|X_{t+i^{\prime}}-X_{t+i}\|_{\alpha}$, for $10\leq i<i^{\prime}\leq d-$ 1. So we just look for coefficients $a_{t,i,j}$ such that $\|\mathcal{T}X_{t+i}\|_{\alpha}=\|X_{t+i}\|_{\alpha}$ and $\|\mathcal{T}X_{t+i^{\prime}}-\mathcal{T}X_{t+i}\|_{\alpha}=\|X_{t+i^{\prime}}-X_{t+i}\|_{\alpha}$. Using the independence and unitary scale of the $Z_{j}$ and Proposition 2, we easily get the expressions $\|\mathcal{T}X_{t+i}\|_{\alpha}^{\alpha}=\sum_{j=0}^{i}|a_{t,i,j}|^{\alpha}$ and $\|\mathcal{T}X_{t+i^{\prime}}-\mathcal{T}X_{t+i}\|_{\alpha}^{\alpha}=\sum_{j=0}^{i^{\prime}}|a_{t,i^{\prime},j}-a_{t,i,j}|^{\alpha}$. Using the stationarity of the increments along with equation (3), we also get $\|X_{t+i^{\prime}}-X_{t+i}\|_{\alpha}^{\alpha}=K_{\alpha,H}^{\alpha}|i^{\prime}-i|^{\alpha H}$ and $\|X_{t+i}\|_{\alpha}^{\alpha}=K_{\alpha,H}^{\alpha}|t+i|^{\alpha H}$, so that we obtain the following system of $d(d+1)/2$ nonlinear equations:
$\left\{\begin{array}[]{ll}(e_{i})&K_{\alpha,H}^{\alpha}|t+i|^{\alpha H}=\sum_{j=0}^{i}|a_{t,i,j}|^{\alpha}\\ (e_{i^{\prime},i})&K_{\alpha,H}^{\alpha}|i^{\prime}-i|^{\alpha H}=\sum_{j=0}^{i^{\prime}}|a_{t,i^{\prime},j}-a_{t,i,j}|^{\alpha},\end{array}\right.$
@@
with equation $(e_{i})$ defined for $i\in\llbracket 0,d-1\rrbracket$ and equation $(e_{i^{\prime},i})$ for $(i,i^{\prime})\in\llbracket 0,d-2\rrbracket\times\llbracket i+1,d-1\rrbracket$. Thanks to an invertible transform of this system, we get the equivalent system of equations displayed in Theorem 1, with $(\mathcal{E}_{i,i})=(e_{i})$ and $(\mathcal{E}_{i^{\prime},i})=(e_{i^{\prime}})-(e_{i^{\prime},i})$.
@@
We now prove the uniqueness of the solution, assuming its existence. First, let’s consider that $H=1/\alpha$. The condition $a_{t,i^{\prime},j}=a_{t,i,j}$ when $j\leq i\leq i^{\prime}$ reduces the problem to the search of $d$ coefficients $a_{t,i,i}$, for $i\in\llbracket 0,d-1\rrbracket$, obtained by the $d$ equations $(\mathcal{E}_{i,i})$, which now write $|a_{t,i,i}|^{\alpha}=K_{\alpha,H}^{\alpha}(t+i)-\sum_{j=0}^{i-1}|a_{t,j,j}|^{\alpha}$. This linear problem can be written with a triangular matrix with nonzero diagonal coefficients, so the solution exists and is unique. Finally, noting that $K_{\alpha,H}=1$ when $H=1/\alpha$, we get the following expression for the coefficients: $a_{t,i,j}=t^{\frac{1}{\alpha}\mathds{1}_{j=0}}\mathds{1}_{i\geq j}$.
@@
We now assume that $H\neq 1/\alpha$. We solve the system iteratively in the lexicographical order of the indices in $(\mathcal{E}_{i^{\prime},i})$. There is no difficulty when $i^{\prime}=i$. When $i^{\prime}\neq i$, we have to prove that the function $f_{t,i^{\prime},i}$ is injective. The function $f_{t,i^{\prime},i}$ is differentiable in $(a_{t,i,i},+\infty)$, which is its domain of definition when $H>1/\alpha$, according to the condition $a_{t,i^{\prime},i}>a_{t,i,i}$ given in the theorem. Its derivative is $f^{\prime}_{t,i^{\prime},i}(z)=\alpha(|z|^{\alpha-1}-|z-a_{t,i,i}|^{\alpha-1})$ in this interval. We note that when $H>1/\alpha$, then $\alpha>1$ since $H<1$. We also note that $|z-a_{t,i,i}|=z-a_{t,i,i}<z$. So $f^{\prime}_{t,i^{\prime},i}$ is strictly positive in $(a_{t,i,i},+\infty)$ and $f_{t,i^{\prime},i}$ is injective.
@@
The function $f_{t,i^{\prime},i}$ is differentiable as well in $(0,a_{t,i,i})$, which is the relevant interval when $H<1/\alpha$, after the condition $a_{t,i^{\prime},i}<a_{t,i,i}$. In this case, $f^{\prime}_{t,i^{\prime},i}(z)=\alpha(|z|^{\alpha-1}+|z-a_{t,i,i}|^{\alpha-1})$, which is again strictly positive in $(0,a_{t,i,i})$, so $f_{t,i^{\prime},i}$ is injective. ∎

### v2

@@ Appendix B Proof of Theorem 1
The autocodifference of the LFSM is provided in equation (5). But the knowledge of the symmetric codifference matrix of $(X_{t},...,X_{t+d-1})^{\prime}$ is equivalent to the knowledge of $\|X_{t+i}\|_{\alpha}$, for $0\leq i\leq d-1$, and $\|X_{t+\ell}-X_{t+i}\|_{\alpha}$, for $0\leq i<\ell\leq d-1$. So we just look for coefficients $a_{t,i,j}$ such that $\|\mathcal{T}X_{t+i}\|_{\alpha}=\|X_{t+i}\|_{\alpha}$ and $\|\mathcal{T}X_{t+\ell}-\mathcal{T}X_{t+i}\|_{\alpha}=\|X_{t+\ell}-X_{t+i}\|_{\alpha}$. Using the independence and unitary scale of the $Z_{j}$ and Proposition 1, we easily get the expressions $\|\mathcal{T}X_{t+i}\|_{\alpha}^{\alpha}=\sum_{j=0}^{i}|a_{t,i,j}|^{\alpha}$ and $\|\mathcal{T}X_{t+\ell}-\mathcal{T}X_{t+i}\|_{\alpha}^{\alpha}=\sum_{j=0}^{\ell}|a_{t,\ell,j}-a_{t,i,j}|^{\alpha}$. Using the stationarity of the increments along with equation (4), we also get $\|X_{t+\ell}-X_{t+i}\|_{\alpha}^{\alpha}=K_{\alpha,H}^{\alpha}|\ell-i|^{\alpha H}$ and $\|X_{t+i}\|_{\alpha}^{\alpha}=K_{\alpha,H}^{\alpha}|t+i|^{\alpha H}$, so that we obtain the following system of $d(d+1)/2$ nonlinear equations:
$\left\{\begin{array}[]{ll}(e_{i})&K_{\alpha,H}^{\alpha}|t+i|^{\alpha H}=\sum_{j=0}^{i}|a_{t,i,j}|^{\alpha}\\ (e_{\ell,i})&K_{\alpha,H}^{\alpha}|\ell-i|^{\alpha H}=\sum_{j=0}^{\ell}|a_{t,\ell,j}-a_{t,i,j}|^{\alpha},\end{array}\right$.
@@
with equation $(e_{i})$ defined for $i\in\llbracket 0,d-1\rrbracket$ and equation $(e_{\ell,i})$ for $(i,\ell)\in\llbracket 0,d-2\rrbracket\times\llbracket i+1,d-1\rrbracket$. Thanks to an invertible transform of this system, we get the equivalent system of equations displayed in Theorem 1, with $(\mathcal{E}_{i,i})=(e_{i})$ and $(\mathcal{E}_{\ell,i})=(e_{\ell})-(e_{\ell,i})$.
@@
We now prove the uniqueness of the solution, assuming its existence. First, let’s consider that $H=1/\alpha$. The condition $a_{t,\ell,j}=a_{t,i,j}$ when $j\leq i\leq\ell$ reduces the problem to the search of $d$ coefficients $a_{t,i,i}$, for $i\in\llbracket 0,d-1\rrbracket$, obtained by the $d$ equations $(\mathcal{E}_{i,i})$, which now write $|a_{t,i,i}|^{\alpha}=K_{\alpha,H}^{\alpha}(t+i)-\sum_{j=0}^{i-1}|a_{t,j,j}|^{\alpha}$. This linear problem can be written with a triangular matrix with nonzero diagonal coefficients, so the solution exists and is unique. Finally, noting that $K_{\alpha,H}=1$ when $H=1/\alpha$, we get the following expression for the coefficients: $a_{t,i,j}=t^{\frac{1}{\alpha}\mathds{1}_{j=0}}\mathds{1}_{i\geq j}$.
@@
We now assume that $H\neq 1/\alpha$. We solve the system iteratively in the lexicographical order of the indices in $(\mathcal{E}_{\ell,i})$. There is no difficulty when $\ell=i$. When $\ell\neq i$, we have to prove that the function $f_{t,i}$ is injective. The function $f_{t,i}$ is differentiable in $(a_{t,i,i},+\infty)$, which is its domain of definition when $H>1/\alpha$, according to the condition $a_{t,\ell,i}>a_{t,i,i}$ given in the theorem. Its derivative is $f^{\prime}_{t,i}(z)=\alpha(|z|^{\alpha-1}-|z-a_{t,i,i}|^{\alpha-1})$ in this interval. We note that when $H>1/\alpha$, then $\alpha>1$ since $H<1$. We also note that $|z-a_{t,i,i}|=z-a_{t,i,i}<z$. So $f^{\prime}_{t,i}$ is strictly positive in $(a_{t,i,i},+\infty)$ and $f_{t,i}$ is injective.
@@
The function $f_{t,i}$ is differentiable as well in $(0,a_{t,i,i})$, which is the relevant interval when $H<1/\alpha$, after the condition $a_{t,\ell,i}<a_{t,i,i}$. In this case, $f^{\prime}_{t,i}(z)=\alpha(|z|^{\alpha-1}+|z-a_{t,i,i}|^{\alpha-1})$, which is again strictly positive in $(0,a_{t,i,i})$, so $f_{t,i}$ is injective. ∎